UT 20120504 AO#2 Commissioning

Observer: SEsposito, RBriguglio, APuglisi, FQuiros-Pacheco (LBTO)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: DMiller
Telescope Support: MMidkiff(LBTO)
Instrument Support: TShih (LBTO)

Summary

Details

01:45 Opened Chamber

02:18 SX AdSec in a strange state. AOSGUI says it is flat, but Alfio says it is not. Alfio fixes it.

02:20 Steve align pointing

02:26 Preset fail. No star.

02:35 Doubled the LUT on the Adsec Control GUI.

02:40 IDL Sext wfsc 10-12 image worse

02:41 Steve Clear C00

02:42 Sext wfsc 14-25 Getting better. Center of subaps not correct.

02:48 Collimated. Click Send Zern to PSF. First application we lost the star because the primary/secondary collimation LUT are applied.

02:50 Steve resend preset. Success.

02:54 Runa sent ACE preset

-- DougMiller - 03 May 2012
Topic revision: r2 - 04 May 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback